OpenStudy (kj4uts):

Shawna and Pranav each improved their yards by planting hostas and geraniums. They bought their supplies from the same store. Shawna spent $130 on 2 hostas and 9 geraniums. Pranav spent $78 on 6 hostas and 1 geranium. What is the cost of one hosta and the cost of one geranium?

Mehek (mehek14):

\(\tt{2h+9g=130\\6h+1g=78}\)

Mehek (mehek14):

do you know how to solve using elimination ?

Mehek (mehek14):

@KJ4UTS ?

OpenStudy (kj4uts):

@Mehek14 sorry about that I had to do something, no I am not sure how to solve using elimination.

Mehek (mehek14):

can you multiply everything in the first equation by 3

Mehek (mehek14):

this will help you eliminate h

OpenStudy (kj4uts):

you mean multiple 2 and 9 by 3?

Mehek (mehek14):

2, 9 , and 130

OpenStudy (kj4uts):

@Mehek14 I was looking at my problem here and it says to solve using substitution is that the same thing as elimination?

Mehek (mehek14):

those are two ways to solve for it you get the same answer using either one

OpenStudy (kj4uts):

oh ok if you don't mind could you explain substitution?

Mehek (mehek14):

Then lets use substitution solve for g in the second equation

OpenStudy (kj4uts):

do I move 6 to the other side 78-6=72 / 1g = 72?

Mehek (mehek14):

\(\tt{g=78-6h}\)

Mehek (mehek14):

not plug the value of g from the second equation for g in the first equation

OpenStudy (kj4uts):

1g + 9g = 10g?

Mehek (mehek14):

\(\tt{2h+9(78-6h)=130}\)

OpenStudy (kj4uts):

h=11

Mehek (mehek14):

good now put 11 for h in any equation

OpenStudy (kj4uts):

2*11=22 130-22=108 / 12 6*11=66 78-66=12 1 hosta $11? 1 geranium $12?

Mehek (mehek14):

correct :)

OpenStudy (kj4uts):

Thank you for your time and help :)

Mehek (mehek14):

yw :)